OpenStudy (anonymous):

If i=square root of -1, what's the value of i^8? Please, I need this one solved; am studying for the COMPASS TEST.

OpenStudy (anonymous):

hello after your dfone helping her could you help me with my questions please? Thanks so much if so.

OpenStudy (anonymous):

Let's group them first: (i*i)*(i*i)*(i*i)*(i*i) then we have -1*-1*-1*-1=-1^4 1*1=1^2 =1

OpenStudy (anonymous):

Thank you, honey. I'm a 39 year old pursuing a bachelor's degree!!

OpenStudy (anonymous):

Good for you - I'm 29 and doing the same. Best of luck in your studies.

OpenStudy (anonymous):

Thank You!!!

OpenStudy (amistre64):

im 39 going for business accounting and a masters in math ;) good luck